Please help to answer the follwing question.
If the profit function for a commodity is p=6400x-18x^2-(1/3)x^3-40,000 dollars, selling how many units, x, will result in a maximum profit? Find the maximum profit.
(I am having troublw getting past this:
p'=6400-36x-x^2=0
Thanks!

Answer by stanbon(75887)

If the profit function for a commodity is p=6400x-18x^2-(1/3)x^3-40,000 dollars, selling how many units, x, will result in a maximum profit? Find the maximum profit.
p'(x) = 6400 - 36x - x^2
----
Solve x^2 + 36x + 6400 = 0
(x+100)(x-64) = 0
Poxitive solution:
x = 64
p(64) = $208,491
----------------------
Cheers,
Stan H.
----------------------